OpenStudy (anonymous):

simplify sqroot 4xy^3 (multiplied by) sqroot 6y

OpenStudy (watchmath):

\(\sqrt{4xy^3}\cdot\sqrt{6y}=\sqrt{6\cdot 4\cdot x\cdot y^4}=2y^2\sqrt{6x}\) Let me know if you have any question.